OpenStudy (anonymous):

Find the general solution of the given differential equation:

OpenStudy (anonymous):

\[y^{(6)}+y=0\]

OpenStudy (anonymous):

I was able to determine the characteristic equation and simplify:\[r^6+1=0 \rightarrow r=\sqrt[6]{-1}\]

OpenStudy (loser66):

No, it is not that. You must derive as \((r^2)^3 +((1)^2)^3=(r^2+1) (r^4-r^2+1)\) and solve for 6 of r

OpenStudy (anonymous):

Oh, I see. How did you know to do it factor it exactly like that?

OpenStudy (loser66):

expand more to get \(-(r^2+1) (-r^2 +\sqrt3 r-1)(r^2 +\sqrt3 r+1)=0\) and manipulate more

OpenStudy (loser66):

You know \(a^3 + b^3\) formula, right?

OpenStudy (anonymous):

Yeah

OpenStudy (loser66):

The leftover is ok, right?

OpenStudy (anonymous):

And then solve the first term algebraically and the last two using quadratic formula, correct?

zepdrix (zepdrix):

This corresponds to the 6th roots of unity :) So if you're comfortable with complex numbers, you could do it that way. Or you can stick with a couple of quadratics as Lose described if that's preferable.

zepdrix (zepdrix):

\[\large\rm r=(\color{orangered}{-1})^{1/6}\]\[\large\rm r=\left(\color{orangered}{e^{i\left(-\frac{\pi}{2}+2k \pi\right)}}\right)^{1/6}\]Which becomes:\[\large\rm r=\color{orangered}{e^{i\left(-\frac{\pi}{12}+\frac{2k \pi}{12}\right)}}\] And then k=0, 1, 2, 3, 4, 5 give you your roots. And then those correspond to a bunch of sines and cosines ya? :o Hmm ya maybe that's more complicated :d

OpenStudy (anonymous):

That's an ugly one, but I think that's the direction I was trying to head toward! ;D I could also use \(\frac{ 3\pi }{ 2 }\) rather than \(-\frac{ \pi }{ 2 }\) no?

zepdrix (zepdrix):

Yes :)

OpenStudy (anonymous):

Mm, yeah and then it will follow the form \(C_1\cos(\theta)+C_2\sin(\theta)\) right? Or should there be an i in front of the sin? I never know when to put the i or not..

zepdrix (zepdrix):

Correct, no i. I'm trying to work through the steps to justify that though. Hehe sec

OpenStudy (anonymous):

Yeah some of the questions had solutions including i with the sin, but most of the time they didn't so I was getting confused DX :P

zepdrix (zepdrix):

hmm

zepdrix (zepdrix):

Oh oh oh, ok. I was getting really confused because Wolfram was giving different angles than what I was coming up with. https://www.wolframalpha.com/input/?i=%28-1%29%5E%281%2F6%29

zepdrix (zepdrix):

Negative 1 is along the real axis, so that's at an angle of pi, not 3pi/2. My bad, my bad.

OpenStudy (anonymous):

Oh I completely missed that mistake too lol

zepdrix (zepdrix):

\[\large\rm r_k^6=\exp\left\{i\left(\pi+2k \pi\right)\right\}\]\[\large\rm r_k=\exp\left\{i\left(\pi+2k \pi\right)\right\}^{1/6}\]\[\large\rm r_k=\exp\left\{i\left(\frac{\pi}{6}+\frac{2k \pi}{6}\right)\right\},\qquad\qquad k=0,1,2,3,4,5\] So for k=0, our first root will be,\[\large\rm r_0=e^{i \pi/6}=\cos\frac{\pi}{6}+i~\sin\frac{\pi}{6}\]\[\large\rm r_0=\frac{\sqrt3}{2}+i~\frac{1}{2}\] Then our first solution: let's call it y_o to match our root notation,\[\large\rm y_0=c_0e^{r_0t}\]\[\large\rm y_0=c_0 ~\exp\left\{\left(\frac{\sqrt3}{2}+i~\frac{1}{2}\right)t\right\}\]\[\large\rm y_0=c_0 e^{\frac{\sqrt{3}}{2}t}\color{orangered}{e^{i\frac{1}{2}t}}\]\[\large\rm y_0=c_0 e^{\frac{\sqrt{3}}{2}t}\color{orangered}{\left(\cos\frac{1}{2}t+i~\sin\frac{1}{2}t\right)}\] Oh yah, don't do it this way >.< Cause then we run into that question with the i again... And I don't wanna have to think about that right now lolol So from here:\[\large\rm r_0=\frac{\sqrt3}{2}+i~\frac{1}{2}\]We get,\[\large\rm y_0=c_0~e^{\frac{\sqrt{3}}{2}t}\cos\frac{t}{2}+c_0~e^{\frac{\sqrt{3}}{2}t}\sin\frac{t}{2}\]Ya? :o

zepdrix (zepdrix):

exp{ } = e^{ } Just in case there was any confusion on that. writing these tiny fractions inside an exponent looks really ugly sometimes.

OpenStudy (anonymous):

That makes sense except what happened to the i in the last step? >_<

zepdrix (zepdrix):

When we have a root: \(\large\rm r=\alpha+\beta i\) Our solution is of the form: \(\large\rm y=c_1 e^{\alpha}\cos\beta+c_2 e^{\alpha}\sin\beta\) I was using that uhhh shortcut thing so we could avoid that question hehe. Where does the i go, that's a good question. I don't think it get's absorbed into the c... so I'm trying to think.. >.<

zepdrix (zepdrix):

Woops*\[\large\rm y=c_1 e^{\alpha t}\cos\beta t+c_2 e^{\alpha t}\sin\beta t\]

OpenStudy (anonymous):

I think it's because the sin of any angle is the vertical component, which is on the imaginary axis. So it already accounts for it perhaps?

OpenStudy (anonymous):

Well not *on* the axis, but in the same direction.

zepdrix (zepdrix):

Ok, I think it's because `complex roots always come in conjugate pairs`. Remember that back from Algebra? So if \(\large\rm r=\frac{\sqrt3}{2}+\frac{1}{2}i\) is one of our roots, Then we also know that \(\large\rm \frac{\sqrt{3}}{2}-\frac{1}{2}i\) is a root! So then a couple of our solutions are,\[\large\rm y=c_1 e^{\sqrt3/2 t+1/2i t}+c_2e^{\sqrt3/2t-1/2i t}\]\[\rm y=c_1 e^{\sqrt3/2 t}\left[\cos\frac{t}{2}+i \sin\frac{t}{2}\right]+c_2e^{\sqrt3/2t}\left[\cos\left(-\frac{t}{2}\right)+i \sin\left(-\frac{t}{2}\right)\right]\]\[\rm y=c_1 e^{\sqrt3/2 t}\left[\cos\frac{t}{2}+i \sin\frac{t}{2}\right]+c_2e^{\sqrt3/2t}\left[\cos\left(\frac{t}{2}\right)-i \sin\left(\frac{t}{2}\right)\right]\]Distribute all of this mess out, Then regroup,\[\rm y= e^{\sqrt3/2 t}\left[(c_1+c_2)\cos\frac{t}{2}+(c_1i-c_2i)\sin\frac{t}{2}\right]\]

zepdrix (zepdrix):

So I guess I was wrong before, we ARE in fact absorbing the i into the c's.

zepdrix (zepdrix):

\[\rm y= e^{\sqrt3/2 t}\left[\color{orangered}{(c_1+c_2)}\cos\frac{t}{2}+\color{orangered}{(c_1i-c_2i)}\sin\frac{t}{2}\right]\]\[\rm y= e^{\sqrt3/2 t}\left[\color{orangered}{c_3}\cos\frac{t}{2}+\color{orangered}{c_4}\sin\frac{t}{2}\right]\]

zepdrix (zepdrix):

Khan explains it in nice detail on youtube if you wanna check that out it some point: https://www.youtube.com/watch?v=6xEO4BeawzA https://www.youtube.com/watch?v=jJyRrIZ595c

zepdrix (zepdrix):

And what you might notice is, I used two of the roots, and ended up with only 1 cosine, and 1 sine. There was overlap, and we were able to combine them together. This should be expected, ya? We shouldn't end up with 12 terms if it's a 6th root, after all we know that we end up with 2 terms (1cosine 1sine) when it's a `square` root.

OpenStudy (anonymous):

Oooooh that makes more sense now. And yeah I was also wondering why we would not have 12 terms because there are cos and sin for each solution. But that makes sense. Thank you!

zepdrix (zepdrix):

cool ୧ʕ•̀ᴥ•́ʔ୨

OpenStudy (anonymous):

d'aww

zepdrix (zepdrix):

XD